I need help with questions 6 - 12! Thank youQuestion 6 2 points Save Answer Suppose the population of Friendly Town on September, 2014 is 45,000. The number of new active cases of gonorrhea for May 1st through September 30th is 34 and the number of active cases on September 30th of gonorrhea is 234. The prevalence rate was 76 per 100,000 population 520 per 100,000 population 596 per 100,000 population 7.6 per 100,000 population 52 per 100,000 population

Question 9

question 10
Answer #1

6. Total population of friendly town =45000

Number of active cases of gonnorrhea =234

Number of new cases of gonnorrhea = 34

Total cases of gonnorrhea = 268

Prevalence rate is the total number of presence of particular disease in a population including both old and new case.

prevalence rate is calculated from 100000 population

So, out of 45000 population 268 has gonnorrhea

Then , for 100000 population it will be 268÷45000 × 100000 =596 /100000 is the prevalence rate

7. Total number of HIV related death under 65 year = 14095 -(509+4)= 13582

14095 = total number of death of all age

509 = number of death above 65year

4 = number of death in which age is not specified

So, mortality rate = number of death below 65 year / number of population below 65 year × 1000

=13582/ 252755 ×1000 = 53.73

8. Standardized mortality rate is the ratio between the number of observed death related to a disease upon number of expected death based on age and sex .

Number of observed death due to cancer =5600

Number of expected death due to cancer =6700

So, standardized mortality rate = 5600/6700 = 0.84

9.HIV related mortality rate for all age group :-

Number of total death of all age group / number of total population including all age group ×1000

= 14095/288357 ×1000

= 48.8